# Anna Fino Dipartimento di Matematica “G. Peano”, Università degli studi di Torino Via Carlo Alberto 10

### v1

@@ 1. Introduction
(Theorem 8). As a consequence, we prove that if $M$ is a compact self-dual simply-connected four-manifold such that its twistor space $Z$ satisfies the $\partial\overline{\partial}$-lemma, then $M$ must be diffeomorphic to either the $4$-sphere or to a connected sum $_{k}\mathbb{C}P^{2}$ for some $k\geq 1$ (Corollary 11). Notably, each of these manifolds admits a self-dual metric whose corresponding twistor space satisfies the $\partial\overline{\partial}$-lemma. Using results from [7], we conclude that the twistor space $Z$ of a compact self-dual simply-connected four-manifold $M$ satisfies the $\partial\overline{\partial}$-lemma if and only if $Z$ lies in Fujiki’s class $\mathcal{C}$, or equivalently, if $Z$ is Moishezon.
@@
Acknowledgments. This work arose from the AIM workshop “Geometric partial differential equations from unified string theories”. The authors are grateful for the great working environment provided by AIM. A. Fino, N. Tardini and A. Tomassini are partially supported by the Project PRIN 2022 “Real and Complex Manifolds: Geometry and Holomorphic Dynamics 2022AP8HZ9” and by GNSAGA of INdAM. G. Grantcharov is partially supported by a grant from the Simons Foundation (#853269). L. Vezzoni is partially supported by the project PRIN 2022 “Differential-geometric aspects of manifolds via Global Analysis” and by GNSAGA of INdAM.
@@ 2.1. Bott-Chern cohomology and $\partial\overline{\partial}$-lemma
Recall that a compact complex manifold is said to satisfy the $\partial\overline{\partial}$-Lemma if the natural vertical map $H^{\bullet,\bullet}_{BC}(X)\longrightarrow H^{\bullet,\bullet}_{A}(X)$ is injective. This is equivalent to any of the above maps being an isomorphism ([9, Lemma 5.15]). By [9] every compact Kähler manifold, and more in general manifolds in class $\mathcal{C}$ of Fujiki (i.e., bimeromorphic to a Kähler manifold), satisfy the $\partial\overline{\partial}$-lemma. Hence the Bott-Chern and Aeppli cohomologies could provide more informations on a compact complex manifold which does not admit any Kähler metric. Moreover, the $\partial\overline{\partial}$-lemma can be numerically characterized in terms of dimensions of cohomology groups as follows.
@@ 2.2. Twistor spaces
In [12, p. 658] Eastwood and Singer, using the Leray-Hirsch theorem, compute the Betti numbers of $Z$:
@@ 3. Bott-Chern and Aeppli cohomologies of twistor spaces and $\partial\overline{\partial}$-lemma
In this section we will describe the Bott-Chern and Aeppli cohomologies of the twistor space of a compact self-dual $4$-manifold and we characterize the validity of the $\partial\overline{\partial}$-lemma. In order to do that we will start with some lemmas.
@@
Let $M$ be a compact self-dual $4$-manifold and let $Z$ be its twistor space. Then, the $(1,0)$ and $(0,1)$ Bott-Chern and Aeppli cohomology groups are as follows
    $H^{1,0}_{BC}(Z)\simeq H^{0,1}_{BC}(Z)=\left\{0\right\}$;
    $H^{1,0}_{A}(Z)\simeq H^{0,1}_{A}(Z)\simeq H^{0,1}_{\overline{\partial}}(Z)$.
@@
We first prove that $H^{1,0}_{BC}(Z)=\left\{0\right\}$. Let $\alpha]\in H^{1,0}_{BC}(Z)=\text{Ker}\partial\cap\text{Ker}\overline{\partial}$, then $\overline{\partial}\alpha=0$, hence $\alpha\in H^{1,0}_{\overline{\partial}}(Z)=\text{Ker}\overline{\partial}=\left\{0\right\}$, hence $\alpha=0$.
We prove now that $H^{1,0}_{A}(Z)\simeq H^{0,1}_{A}(Z)\simeq H^{0,1}_{\overline{\partial}}(Z)$. In fact the map
$H^{0,1}_{A}(Z)\to H^{0,1}_{\overline{\partial}}(Z),\quad[\alpha]\mapsto[\alpha]$
is well-defined. Indeed, if $\alpha]\in H^{0,1}_{A}(Z)$ then $\partial\overline{\partial}\alpha=0$ hence $\overline{\partial}\alpha$ is a $\partial$-closed $(0,2)$-form, but $H^{0,2}_{\partial}(Z)\simeq H^{2,0}_{\overline{\partial}}(Z)=\left\{0\right\}$, hence $\overline{\partial}\alpha=0$ and so $\alpha$ defines a class in $H^{0,1}_{\overline{\partial}}(Z)$. We only need to show that this map is injective. Now let $\alpha]\in H^{0,1}_{A}(Z)$ such that $\alpha]=0\in H^{0,1}_{\overline{\partial}}(Z)$, namely $\alpha=\overline{\partial}f$, but this means that $\alpha]=0\in H^{0,1}_{A}(Z)$, concluding the proof. ∎
@@
Let $M$ be a compact self-dual $4$-manifold and let $Z$ be its twistor space. Then, the $(2,0)$ and $(0,2)$ Bott-Chern and Aeppli cohomology groups are as follows
    $H^{2,0}_{BC}(Z)\simeq H^{0,2}_{BC}(Z)=\left\{0\right\}$;
    $H^{2,0}_{A}(Z)\simeq H^{0,2}_{A}(Z)\simeq H^{0,2}_{\overline{\partial}}(Z)$.
@@
We first prove that $H^{2,0}_{BC}(Z)=\left\{0\right\}$. Let $\alpha]\in H^{2,0}_{BC}(Z)=\text{Ker}\partial\cap\text{Ker}\overline{\partial}$, then $\overline{\partial}\alpha=0$, hence $\alpha\in H^{2,0}_{\overline{\partial}}(Z)=\text{Ker}\overline{\partial}=\left\{0\right\}$, hence $\alpha=0$.
We prove now that $H^{2,0}_{A}(Z)\simeq H^{0,2}_{A}(Z)\simeq H^{0,2}_{\overline{\partial}}(Z)$. In fact the map
$H^{0,2}_{A}(Z)\to H^{0,2}_{\overline{\partial}}(Z),\quad[\alpha]\mapsto[\alpha]$
is well-defined. Indeed, if $\alpha]\in H^{0,2}_{A}(Z)$ then $\partial\overline{\partial}\alpha=0$ hence $\overline{\partial}\alpha$ is a $\partial$-closed $(0,3)$-form, but $H^{0,3}_{\partial}(Z)\simeq H^{3,0}_{\overline{\partial}}(Z)=\left\{0\right\}$, hence $\overline{\partial}\alpha=0$ and so $\alpha$ defines a class in $H^{0,2}_{\overline{\partial}}(Z)$. We only need to show that this map is injective. Now let $\alpha]\in H^{0,2}_{A}(Z)$ such that $\alpha]=0\in H^{0,2}_{\overline{\partial}}(Z)$, namely $\alpha=\overline{\partial}\beta$, but this means that $\alpha]=0\in H^{0,2}_{A}(Z)$, concluding the proof. ∎
@@
Let $M$ be a compact self-dual $4$-manifold and let $Z$ be its twistor space. Then, the $(3,0)$ and $(0,3)$ Bott-Chern and Aeppli cohomology groups are as follows
    $H^{3,0}_{BC}(Z)\simeq H^{0,3}_{BC}(Z)=\left\{0\right\}$;
    $H^{3,0}_{A}(Z)\simeq H^{0,3}_{A}(Z)=\left\{0\right\}$.
@@
We first prove that $H^{3,0}_{BC}(Z)=\left\{0\right\}$. Let $\alpha]\in H^{3,0}_{BC}(Z)=\text{Ker}\partial\cap\text{Ker}\overline{\partial}$, then $\overline{\partial}\alpha=0$, hence $\alpha\in H^{3,0}_{\overline{\partial}}(Z)=\text{Ker}\overline{\partial}=\left\{0\right\}$, hence $\alpha=0$.
The fact that $H^{3,0}_{A}(Z)\simeq H^{0,3}_{A}(Z)=\left\{0\right\}$ follows by duality between the Bott-Chern and Aeppli cohomology groups. ∎
@@
The second equality $h^{2,0}_{A}(Z)=h^{0,2}_{A}(Z)=b_{-}(M)$ follows from Lemma 3 saying that
@@
In case $M$ is simply connected we can characterize manifolds whose twistor space satisfies the $\partial\overline{\partial}$-lemma.
Corollary 11.
Let $M$ be a compact self-dual simply-connected 4-manifold such that its twistor space $Z$ satisfies $\partial\overline{\partial}$-lemma. Then $M$ is diffeomorphic to the 4-sphere or a connected sum $\sharp_{k}\mathbb{C}P^{2}$ of $k$ copies of $\mathbb{C}P^{2}$ for $k=1,2,\ldots\,$. Each of these spaces admits a self-dual metric for which $Z$ satisfies $\partial\overline{\partial}$-lemma.
Proof.
Observe that, under the assumption, $M$ has $b_{1}(M)=b_{-}(M)=0$ and $\pi_{1}(M)=0$ and so using Donaldson’s Theorem [10, 13] we obtain that $M$ is diffeomorphic to the 4-sphere or a connected sum $\sharp_{k}\mathbb{C}P^{2}$ of $k$ copies of $\mathbb{C}P^{2}$ for $k=1,2,\ldots\,$. For the last statement we refer to [15] and [17]. ∎
As a consequence, from [7] it follows
Corollary 12.
Let $M$ be a compact self-dual simply-connected 4-manifold and denote with $Z$ its twistor space. Then, the following are equivalent:
  • $Z$ satisfies $\partial\overline{\partial}$-lemma;
@@
By [12, Corollary 4.5] since $M$ is Einstein it is regular (in the sense of [12, page 661]) and so we can apply [12, Theorem 4.6]. More precisely, since $M$ is regular and $E_{1}=E_{\infty}$, by [12, Theorem 4.6] there is a short exact sequence
@@
However, by [12, page 658] the kernel of the map $H^{1,1}_{\overline{\partial}}(Z)\to H^{2,1}_{\overline{\partial}}(Z)$ is canonically identified with $\mathbb{C}\oplus H^{2}_{-}(M,\mathbb{C})$ and so it has dimension $b_{-}+1=1$. But this is absurd since, by $h^{2,1}_{\overline{\partial}}(Z)=0$ such map is trivial $H^{1,1}_{\overline{\partial}}(Z)\to H^{2,1}_{\overline{\partial}}(Z)=\left\{0\right\}$ and the kernel should have dimension $h^{1,1}_{\overline{\partial}}(Z)=2$.
@@
Recall that by [12, Corollary 5.7], for any $g\geq 0$, there are metrics on $M=\sharp_{g}(\mathbb{S}^{1}\times\mathbb{S}^{3})$ such that the Frölicher spectral sequence of the associated twistor space degenerates at the first step. We can therefore compute the Hodge numbers of the twistor space using (2). Indeed, $h^{0,1}_{\overline{\partial}}=b_{1}=1$,
@@ 4. Dolbeault cohomology of the twistor space of the torus
In this section we consider the twistor space $Z$ of the flat $4$-dimensional torus, which is known to not satisfy the $\partial\overline{\partial}$ lemma, and we compute explicitly the Dolbeault cohomology. The calculations could be generalized to higher-dimensional tori. The notations follow [14].
The twistor space $Z$ is the quotient of the bundle ${\mathfrak{t}}^{1,0}\otimes{\it O}(1)$ on $\mathbb{C}P^{1}\equiv\mathbb{S}^{2}$ where ${\mathfrak{t}}^{1,0}$ is the $(1,0)$-part of the complexification of the $4$-dimensional Abelian algebra $\mathfrak{t}$ ([5, Ex. 13.64, 13.66])
Let $\lambda_{1},\lambda_{2}]$ be the homogeneous coordinates on $\mathbb{C}P^{1}$. On $U_{1}=\{\lambda\in\mathbb{C}P^{1}:\lambda_{1}\neq 0\}$, define $\nu=\frac{\lambda_{2}}{\lambda_{1}}$. On $U_{2}=\{\lambda\in\mathbb{C}P^{1}:\lambda_{2}\neq 0\}$, define $\mu=\frac{\lambda_{1}}{\lambda_{2}}$.
On $U_{2}$, the (0,1)-forms
(3) ${\overline{\sigma}}_{1}=\frac{\overline{\mu}d\overline{z_{1}}-d{z}_{2}}{1+|\mu|^{2}},\hskip 14.45377pt{\overline{\sigma}}_{2}=\frac{\overline{\mu}d\overline{z_{2}}+d{z}_{1}}{1+|\mu|^{2}}$
are holomorphic because
@@
According to [14] on $Z$ one has
$H^{k}(Z,p^{*}{\it O}(\ell))=\mathfrak{t}^{*(0,k)}\otimes S^{\ell+k}\mathbb{C}^{2}$
for $k=0,1,2$ and the corresponding Dolbeault groups $H^{0,k}(Z)$ can be explicitly described. Define
@@
Then $\{\lambda_{1}\overline{\Omega}_{1},\lambda_{2}\overline{\Omega}_{1},\lambda_{1}\overline{\Omega}_{2},\lambda_{2}\overline{\Omega}_{2}\}$ form a basis for the space $H^{1}(Z,{\it O})$. Also $H^{2}(Z,{\it O})$, is spanned by the twisted 2-forms
@@
The space $H^{1}(Z,{\it O})$ has an alternative description. For $k=0,1,2,3$ and over $p^{-1}({\vec{a}})$, define 1-forms
@@
where $I_{0}=Id$, and $I_{1},I_{2},I_{3}$ form the standard hypercomplex structure, so $dz_{1}=dx+iI_{1}dx$ and $dz_{2}=I_{2}dx=iI_{3}dx$, and $I_{\vec{a}}=a_{1}I_{1}+a_{2}I_{2}+a_{3}I_{3}$ for $\vec{a}=(a_{1}.a_{2}.a_{3})\in S^{2}$. Note that the norms of $\omega_{k}$ are all equal to $\sqrt{2}$. Then the space $H^{1}(Z,{\it O})$ is spanned by $\overline{\omega}_{k}$ because
@@
Moreover, again by [12, Theorem 5.3], the Frölicher spectral sequence degenerates of $Z$ at the first step $E_{1}=E_{\infty}$ hence

### v2

@@ 1. Introduction
(Theorem 8). As a consequence, we prove that if $M$ is a compact self-dual simply-connected four-manifold such that its twistor space $Z$ satisfies the $\partial\overline{\partial}$-lemma, then $M$ must be diffeomorphic to either the $4$-sphere or to a connected sum $_{k}\mathbb{C}P^{2}$ for some $k\geq 1$. Notably, each of these manifolds admits a self-dual metric whose corresponding twistor space satisfies the $\partial\overline{\partial}$-lemma. Using results from [7] and [8], we conclude that the twistor space $Z$ of a compact self-dual simply-connected four-manifold $M$ satisfies the $\partial\overline{\partial}$-lemma if and only if $Z$ lies in Fujiki’s class $\mathcal{C}$, or equivalently, if $Z$ is Moishezon.
@@
Acknowledgments. The authors are grateful to the anonymous reviewer for the insightful comments that improved the presentation of the paper. This work arose from the AIM workshop “Geometric partial differential equations from unified string theories”. The authors are grateful for the great working environment provided by AIM. A. Fino, N. Tardini and A. Tomassini are partially supported by the Project PRIN 2022 “Real and Complex Manifolds: Geometry and Holomorphic Dynamics 2022AP8HZ9” and by GNSAGA of INdAM. G. Grantcharov is partially supported by a grant from the Simons Foundation (#853269). L. Vezzoni is partially supported by the project PRIN 2022 “Differential-geometric aspects of manifolds via Global Analysis” and by GNSAGA of INdAM. We also thank M. Verbitsky for pointing out to us the reference [8] and L. Di Cerbo for the information on fake projective planes.
@@ 2.1. Bott-Chern cohomology and $\partial\overline{\partial}$-lemma
Recall that a compact complex manifold is said to satisfy the $\partial\overline{\partial}$-Lemma if the natural vertical map $H^{\bullet,\bullet}_{BC}(X)\longrightarrow H^{\bullet,\bullet}_{A}(X)$ is injective. This is equivalent to any of the above maps being an isomorphism ([10, Lemma 5.15]). By [10] every compact Kähler manifold, and more in general manifolds in class $\mathcal{C}$ of Fujiki (i.e., bimeromorphic to a Kähler manifold), satisfy the $\partial\overline{\partial}$-lemma. Hence the Bott-Chern and Aeppli cohomologies could provide more informations on a compact complex manifold which does not admit any Kähler metric. Moreover, the $\partial\overline{\partial}$-lemma can be numerically characterized in terms of dimensions of cohomology groups as follows.
@@ 2.2. Twistor spaces
In [13, p. 658] Eastwood and Singer, using the Leray-Hirsch theorem, compute the Betti numbers of $Z$:
@@ 3. Bott-Chern and Aeppli cohomologies of twistor spaces and $\partial\overline{\partial}$-lemma
In this section we will describe the Bott-Chern and Aeppli cohomologies of the twistor space of a compact self-dual $4$-manifold and we characterize the validity of the $\partial\overline{\partial}$-lemma. In order to do that we will start with a general lemma.
@@
Let $X$ be a compact complex manifold. Then, the following sequences are exact:
    $0\to H^{p,0}_{BC}(X)\to H^{p,0}_{\overline{\partial}}(X)$;
    $0\to H^{0,q}_{\overline{\partial}}(X)\to H^{0,q}_{A}(X)\xrightarrow{\overline{\partial}}H^{0,q+1}_{\partial}(X)$
@@
We first want to show that for any $p$, the map $H^{p,0}_{BC}(X)\to H^{p,0}_{\overline{\partial}}(X)$ is injective. Notice that $H^{p,0}_{BC}(X)=\text{Ker}\,\partial\cap\text{Ker}\,\overline{\partial}\cap A^{p,0}(X)$ and $H^{p,0}_{\overline{\partial}}(X)=\text{Ker}\,\overline{\partial}\cap A^{p,0}(X)$. The injectivity is then straightforward. Now we show that the map $\iota:H^{0,q}_{\overline{\partial}}(X)\to H^{0,q}_{A}(X)$ is injective for any $q$. Indeed, let $\alpha]_{\overline{\partial}}\in H^{0,q}_{\overline{\partial}}(X)$ such that $\iota[\alpha]_{\overline{\partial}}=[\alpha]_{A}=0$. Since $H^{0,q}_{A}(X)=\frac{\text{Ker}\,\partial\overline{\partial}}{\text{Im}\,\overline{\partial}}$, then $\alpha$ is $\overline{\partial}$-exact and so $\alpha]_{\overline{\partial}}=0$. Finally, consider the map $H^{0,q}_{A}(X)\xrightarrow{\overline{\partial}}H^{0,q+1}_{\partial}(X)=\text{Ker}\,\partial\cap A^{0,q+1}(X)$, defined by $\overline{\partial}[\alpha]_{A}=[\overline{\partial}\alpha]_{\partial}$. Then it is clear that $\text{Im}\,\iota=\text{Ker}\,\overline{\partial}$. ∎
Now we can use this lemma to obtain the following
@@
Let $M$ be a compact self-dual $4$-manifold and let $Z$ be its twistor space. Then, for $p=1,2,3$, the $(p,0)$ and $(0,p)$ Bott-Chern cohomology groups are as follows
$H^{p,0}_{BC}(Z)=H^{0,p}_{BC}(Z)=\left\{0\right\}\,.$
@@
Since by (2), $H^{p,0}_{\overline{\partial}}(Z)=\left\{0\right\}$ for $p=1,2,3$, it follows by Lemma 2 that $H^{p,0}_{BC}(Z)\simeq H^{0,p}_{BC}(Z)=\left\{0\right\}\,$. ∎
@@
Let $M$ be a compact self-dual $4$-manifold and let $Z$ be its twistor space. Then, for $q=1,2,3$, the $(q,0)$ and $(0,q)$ Aeppli cohomology groups are as follows
$H^{q,0}_{A}(Z)=H^{0,q}_{A}(Z)\simeq H^{0,q}_{\overline{\partial}}(Z)\,.$
In particular, $H^{3,0}_{A}(Z)=H^{0,3}_{A}(Z)=\left\{0\right\}$.
@@
Notice that by (2), for $r=0,1,2,3$, $H^{0,r+1}_{\partial}(Z)=\overline{H^{r+1,0}_{\overline{\partial}}(Z)}=\left\{0\right\}$. Hence, applying Lemma 2 we obtain isomorphisms $H^{q,0}_{A}(Z)=H^{0,q}_{A}(Z)\simeq H^{0,q}_{\overline{\partial}}(Z)$
@@
The second equality $h^{2,0}_{A}(Z)=h^{0,2}_{A}(Z)=b_{-}(M)$ again follows from Lemma 4 saying that
@@
Recall that if $M$ is a compact self-dual simply-connected 4-manifold and $Z$ is its twistor space, then by [7] and [8] the following are equivalent:
  • (i)
@@
In particular, $Z$ satisfies the $\partial\overline{\partial}$-lemma. However notice that there are small deformations of $Z$ which do not satisfy (i) and (ii), but they do satisfy the $\partial\overline{\partial}$-lemma. Note that a compact self-dual simply-connected 4-manifold is homeomorphic to $S^{4}$ or a connected sum of $k$ copies, $k\geq 1$ of complex projective planes by [11].
@@
By [13, Corollary 4.5] since $M$ is Einstein it is regular (in the sense of [13, page 661]) and so we can apply [13, Theorem 4.6]. More precisely, since $M$ is regular and $E_{1}=E_{\infty}$, by [13, Theorem 4.6] there is a short exact sequence
@@
However, by [13, page 658] the kernel of the map $H^{1,1}_{\overline{\partial}}(Z)\to H^{2,1}_{\overline{\partial}}(Z)$ is canonically identified with $\mathbb{C}\oplus H^{2}_{-}(M,\mathbb{C})$ and so it has dimension $b_{-}+1=1$. But this is absurd since, by $h^{2,1}_{\overline{\partial}}(Z)=0$ such map is trivial $H^{1,1}_{\overline{\partial}}(Z)\to H^{2,1}_{\overline{\partial}}(Z)=\left\{0\right\}$ and the kernel should have dimension $h^{1,1}_{\overline{\partial}}(Z)=2$.
@@
Recall that by [13, Corollary 5.7], for any $g\geq 0$, there are metrics on $M=\sharp_{g}(\mathbb{S}^{1}\times\mathbb{S}^{3})$ such that the Frölicher spectral sequence of the associated twistor space degenerates at the first step. We can therefore compute the Hodge numbers of the twistor space using (2). Indeed, $h^{0,1}_{\overline{\partial}}=b_{1}=1$,
@@ 4. Dolbeault cohomology of the twistor space of the torus
In this section we consider the twistor space $Z$ of the flat $4$-dimensional torus, which is known to not satisfy the $\partial\overline{\partial}$ lemma, and we compute explicitly the Dolbeault cohomology. The calculations could be generalized to higher-dimensional tori.
The twistor space $Z$ is the quotient of the bundle ${\mathfrak{t}}^{1,0}\otimes{\it O}(1)$ on $\mathbb{C}P^{1}\equiv\mathbb{S}^{2}$ where ${\mathfrak{t}}^{1,0}$ is the $(1,0)$-part of the complexification of the $4$-dimensional Abelian algebra $\mathfrak{t}$ ([5, Ex. 13.64, 13.66]). More precisely, let $\lambda_{1},\lambda_{2}]$ be the homogeneous coordinates on $\mathbb{C}P^{1}$. On $U_{1}=\{\lambda\in\mathbb{C}P^{1}:\lambda_{1}\neq 0\}$, define $\nu=\frac{\lambda_{2}}{\lambda_{1}}$. On $U_{2}=\{\lambda\in\mathbb{C}P^{1}:\lambda_{2}\neq 0\}$, define $\mu=\frac{\lambda_{1}}{\lambda_{2}}$.
As described in [5], on $U_{2}$, the space of (0,1)-forms
(3) ${\overline{\sigma}}_{1}=\frac{\overline{\mu}d\overline{z_{1}}-d{z}_{2}}{1+|\mu|^{2}},\hskip 14.45377pt{\overline{\sigma}}_{2}=\frac{\overline{\mu}d\overline{z_{2}}+d{z}_{1}}{1+|\mu|^{2}},\hskip 14.45377pt{\overline{d}\mu}$
spans $T^{(0,1)}Z$ at each point. Moreover, the forms $\overline{\sigma}_{1},\overline{\sigma}_{2}$} are holomorphic because
@@
As described in [15] on $Z$ one has
$H^{k}(Z,p^{*}{\mathcal{O}}(\ell))=\mathfrak{t}^{*(0,k)}\otimes S^{\ell+k}\mathbb{C}^{2}$
for $k=0,1,2$, where $p$ is the projection from $Z$ to $\mathbb{CP}^{1}$, and the corresponding Dolbeault groups $H^{k}(Z,{\mathcal{O}})=H^{0,k}_{\overline{\partial}}(Z)$ can be explicitly described. Define
@@
Then $\{\lambda_{1}\overline{\Omega}_{1},\lambda_{2}\overline{\Omega}_{1},\lambda_{1}\overline{\Omega}_{2},\lambda_{2}\overline{\Omega}_{2}\}$ form a basis for the space $H^{1}(Z,{\mathcal{O}})=H^{0,1}_{\overline{\partial}}(Z)$. Also $H^{2}(Z,{\mathcal{O}})=H^{0,1}_{\overline{\partial}}(Z)$, is spanned by the twisted 2-forms
@@
The space $H^{1}(Z,{\mathcal{O}})=H^{0,1}_{\overline{\partial}}(Z)$ has an alternative description. For $k=0,1,2,3$ and over $p^{-1}({\vec{a}})$, define 1-forms
@@
where $I_{0}=Id$, and $I_{1},I_{2},I_{3}$ form the standard hypercomplex structure, so $dz_{1}=dx+iI_{1}dx$ and $dz_{2}=I_{2}dx=iI_{3}dx$, and $I_{\vec{a}}=a_{1}I_{1}+a_{2}I_{2}+a_{3}I_{3}$ for $\vec{a}=(a_{1}.a_{2}.a_{3})\in S^{2}$. Note that the norms of $\omega_{k}$ are all equal to $\sqrt{2}$. Then the space $H^{0,1}_{\overline{\partial}}(Z)$ is spanned by $\overline{\omega}_{k}$ because
@@
Moreover, again by [13, Theorem 5.3], the Frölicher spectral sequence degenerates of $Z$ at the first step $E_{1}=E_{\infty}$ hence
@@
    For explanatory purposes, we will compute $\overline{\partial}\left(\bar{\lambda}_{1}\Omega_{1}\wedge\lambda_{1}\bar{\Omega}_{2}+\bar{\lambda}_{2}\Omega_{1}\wedge\lambda_{2}\bar{\Omega}_{2}\right)$ recalling that $\lambda_{1}\bar{\Omega}_{2}$ and $\lambda_{2}\bar{\Omega}_{2}$ are $\overline{\partial}$-closed. Using formulas (9) we have to compute $\overline{\partial}\left(\bar{\mu}\sigma_{1}\wedge\mu\bar{\sigma}_{2}+\sigma_{1}\wedge\bar{\sigma}_{2}\right)$. $\begin{aligned} \overline{\partial}\left(\bar{\mu}\sigma_{1}\wedge\mu\bar{\sigma}_{2}+\sigma_{1}\wedge\bar{\sigma}_{2}\right)=&d\bar{\mu}\wedge\sigma_{1}\wedge\mu\bar{\sigma}_{2}+\frac{\bar{\mu}}{1+|\mu|^{2}}\left(d\mu\wedge\bar{\sigma}_{2}-\mu d\bar{\mu}\wedge\sigma_{1}\right)\wedge\mu\bar{\sigma}_{2}+\\ &+\frac{1}{1+|\mu|^{2}}\left(d\mu\wedge\bar{\sigma}_{2}-\mu d\bar{\mu}\wedge\sigma_{1}\right)\wedge\bar{\sigma}_{2}=\\ =&\left(\mu-\frac{|\mu|^{2}\mu}{1+|\mu|^{2}}\right)d\bar{\mu}\wedge\sigma_{1}\wedge\bar{\sigma}_{2}-\frac{\mu}{1+|\mu|^{2}}d\bar{\mu}\wedge\sigma_{1}\wedge\bar{\sigma}_{2}=0\,.\end{aligned}$
    The differential of the other forms can be computed similarly.
@@
    For explanatory purposes, we will show that $\bar{\lambda}_{1}\Omega_{1}\wedge\lambda_{1}\bar{\Omega}_{1}\wedge\lambda_{1}\bar{\Omega}_{2}+\bar{\lambda}_{2}\Omega_{1}\wedge\lambda_{1}\bar{\Omega}_{1}\wedge\lambda_{2}\bar{\Omega}_{2}$ is $\overline{\partial}$-closed. Indeed, notice that
    $\bar{\lambda}_{1}\Omega_{1}\wedge\lambda_{1}\bar{\Omega}_{1}\wedge\lambda_{1}\bar{\Omega}_{2}+\bar{\lambda}_{2}\Omega_{1}\wedge\lambda_{1}\bar{\Omega}_{1}\wedge\lambda_{2}\bar{\Omega}_{2}=-\left(\bar{\lambda}_{1}\Omega_{1}\wedge\lambda_{1}\bar{\Omega}_{2}+\bar{\lambda}_{2}\Omega_{1}\wedge\lambda_{2}\bar{\Omega}_{2}\right)\wedge\lambda_{1}\bar{\Omega}_{1}$
    and by the computations of $H^{1,1}_{\overline{\partial}}(Z)$ and $H^{0,1}_{\overline{\partial}}(Z)$ we have that both $\bar{\lambda}_{1}\Omega_{1}\wedge\lambda_{1}\bar{\Omega}_{2}+\bar{\lambda}_{2}\Omega_{1}\wedge\lambda_{2}\bar{\Omega}_{2}$ and $\lambda_{1}\bar{\Omega}_{1}$ are $\overline{\partial}$-closed.
    The differential of the other forms can be computed similarly.